OpenStudy (anonymous):

Starting from rest, an object slides freely down an inclined plane in 2.0 seconds with an acceleration of 6.0 meters per second per second. What is the length of the plane? A. 20 m B. 12 m C. 6.0 m D. 2.0 m

OpenStudy (anonymous):

Apply One dimensional Motion equation Find Final velocity = V = U +at V = 6 * 2 V= 12ms V^2 = U^2 + 2as 12^2 = 2* 8 *s S = ?